OpenStudy (anonymous):

Prove 2sinx/sin(2x)+tanx/sinx=2secx

OpenStudy (anonymous):

Use \(\Large \tan x = \frac{\sin x}{\cos x}\) and \(\sin 2x = 2 \sin x \cos x\)

OpenStudy (anonymous):

2sinx/sin2x +tanx/sinx= 2sinx/2sinx.cosx +sinx/cosx/sinx= secx+secx= 2secx